Can someone please help me?I haven't been able to understand how to do this problems very well..
The number of dollars per month it costs you to own a car is a function of the number of kilometers per month you drive it. Based on the information in an issue of time magazine, the costs varies linearly with the distance , and is $366 per month for 300 km per month, and $510 per month for 1500 km per month.
A)write the particular equation expressing cost in terms of distance.
B) About how far could you drive in a month without exceeding a monthly cost of $600?
C) What does the slope represent?
D) List all the reasons you can think of to explain why the dollars per month intercept is grater than zero.
E)Predict the monthly cost if you drive 500, 1000, 2000 miles per month.

Answer by ewatrrr(24785)

 
Hi,
costs varies linearly with the distance
(1500,510)
(300,366) m = (510-366)/(1500-300) = 144/1200 = 12/100 = .12
y = .12x + b
510 = .12*1500 + b
330 = b
f(x) = .12x + 330
.12x + 330 <$600
.12x < 270
x < 2250mi
slope represents the constant of increase cost/mi
b = $330 relates to the basic cost of owning a vehicle
f(500) = $390
f(1000) = $450
f(2000) = $570